OpenStudy (anonymous):

GIVING MEDAL AND BE A FAN! For f(x)=3x+1 g(x)=x^2-6, find (f-g)(x).

jimthompson5910 (jim_thompson5910):

(f-g)(x) is the same as f(x) - g(x)

jimthompson5910 (jim_thompson5910):

are you able to find f(x) - g(x) ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

f(x) - g(x) = (3x + 1) - (x^2 - 6) f(x) - g(x) = 3x + 1 - x^2 + 6 ... distribute f(x) - g(x) = ???

OpenStudy (anonymous):

I honestly don't understand how to work it out at all

OpenStudy (anonymous):

uhhhhh...

jimthompson5910 (jim_thompson5910):

Simplify 3x + 1 - x^2 + 6 to get your answer

OpenStudy (anonymous):

2x^2 +7??????????

jimthompson5910 (jim_thompson5910):

you have 1+6 = 7 correct, but the other terms do not combine (since they aren't like terms)

jimthompson5910 (jim_thompson5910):

so the answer is just \(\Large f(x) - g(x) = -x^2 + 3x + 7\)

OpenStudy (anonymous):

ohhh so you just leave those terms alone?

jimthompson5910 (jim_thompson5910):

yep, if you had -x + 3x then you could combine them to get 2x (since they both are like terms) but in this case, they aren't like terms, so you just leave them as they are

OpenStudy (anonymous):

Are you sure its not -x^2+3x-5?

jimthompson5910 (jim_thompson5910):

yes, that is if f(x)=3x+1 and g(x)=x^2-6

jimthompson5910 (jim_thompson5910):

it would be -x^2+3x-5 if you wanted to find f(x) + g(x)